# Preference-Guided Diffusion for Multi-Objective Offline Optimization

### v1

@@ 4.1 Preference Guided Diffusion
In this work, we explore an alternate guidance strategy that does not involve training surrogate models for every objective. Instead, we train a preference model that predicts whether a design Pareto dominates (Definition 2.1) another design. Given two designs ${\bm{x}}$ and $\hat{\bm{x}}$, we train a (time-conditioned) binary classifier that predicts $p_{\phi}({\bm{x}}\prec\hat{\bm{x}}\mid{\bm{x}},\hat{\bm{x}},t)$. We parameterize this distribution with a multi-layer perceptron (MLP) that takes in two inputs (designs) of size $2\times d$ and outputs the logit of the Bernoulli distribution predicting whether the first input Pareto dominates the second input.
@@
To train the preference classifier, we first sort the points in the training data by their Pareto dominance. The data is then divided into multiple fronts in increasing order of dominance, with points in the same front considered equally dominant. Next, We select $({\bm{x}},\hat{\bm{x}})$ pairs randomly from the dataset. If one design strictly dominates the other, it is labeled as preferred. Otherwise, if there is no strict dominance between either of the selected designs $({\bm{x}},\hat{\bm{x}})$, we assign the label ${\bm{x}}\prec\hat{\bm{x}}$ if ${\bm{x}}$ has more diversity contribution than $\hat{\bm{x}}$ wrt the other designs in the dataset that belong to the same Pareto front (and vice versa). In this work, we calculate the diversity contribution of each point by computing the crowding distance [10] of a selected design w.r.t all other points that belong to the same front in the dataset. Crowding distance for any point ${\bm{x}}$ is computed as follows:
$d_{\text{CD}}({\bm{x}})=\sum_{i=1}^{m}\frac{y_{i}^{+}-y_{i}^{-}}{y_{i}^{{\mathcal{D}}(\text{max})}-y_{i}^{{\mathcal{D}}(\text{min})}},$ (7)
where $y_{i}^{+}$ and $y_{i}^{-}$ are the $i$th objective values of neighboring designs of ${\bm{x}}$ in the corresponding front sorted according to $i$th objective value. $f_{i}^{{\mathcal{D}}(\text{max})}$ and $f_{i}^{{\mathcal{D}}(\text{min})}$ are maximum and minimum values of objective $i$ in the entire dataset. Crowding distance has been used as a secondary selection criterion in evolutionary algorithms like NSGA-2 [10] to maintain diversity of solutions. Using crowding distance to create a binary label encourages the preference model to not only guide the diffusion model towards more Pareto-dominant regions but also ensure that the designs that make up the Pareto front are diverse.
For the denoising model, we train an unconditional diffusion model $\epsilon_{\theta}({\bm{x}}_{t},t)$ (§ 2.2) on the designs ${\bm{x}}$ in the dataset, similar to DDPM [14].
@@
1:  $r\leftarrow{\bm{x}}^{{\mathcal{D}}(\text{best})}$
2:  $\tilde{{\bm{x}}}_{T}\sim{\mathcal{N}}(0,{\mathbb{I}}_{d})$
3:  for t = $T$ to $1$ do
@@
With a trained denoising model $\epsilon_{\theta}({\bm{x}}_{t},t)$ and preference model $p_{\phi}({\bm{x}}\prec\hat{\bm{x}}\mid{\bm{x}},\hat{\bm{x}},t)$, we sample a new design by using classifier guidance (§ 2.3). We input both the denoised variable at the current timestep $\tilde{\bm{x}}_{t}$ as well as from the previous realization of denoising, i.e., $\tilde{\bm{x}}_{t+1}$ for the preference model to estimate $\nabla_{\tilde{\bm{x}}_{t}}p_{\phi}(\tilde{\bm{x}}_{t}\prec\tilde{\bm{x}}_{t+1}\mid\tilde{\bm{x}}_{t},\tilde{\bm{x}}_{t+1},t)$. The sampling procedure is summarized in Algorithm 1. Intuitively, a preference model that generalizes well beyond its training data should guide the denoising process such that, at each step of denoising, the resulting sample $\tilde{\bm{x}}_{t}$ is in a more Pareto-dominant region. With enough timesteps, the resulting denoised sample $\tilde{\bm{x}}_{0}$ will be close to the Pareto front. This approach does not require training surrogate models, thus providing a simple alternative approach to offline MOO. We find that the preference model generalizes well outside of the training data (see Fig. 1), therefore providing guidance to the diffusion model to generate designs outside of the training data close to the Pareto front, while maintaining diversity in the samples.
@@ 5 Experiments
Our evaluation closely follows the benchmarking effort provided in prior work  [37]. We evaluate our approach on two sets of tasks - synthetic and real-world applications-based RE engineering suite [32].
  1. 1.
    Synthetic tasks consist of several subtasks wherein the objective functions are hand-designed with known ground truth Pareto-fronts. These subtasks have been widely used in MOO problems to study the performance of the algorithm. These subtasks consist of 2-3 objectives with $d$ ranging from 10 to 30. We use the same dataset as prior work [37], which consists of 60,000 offline data points.
  2. 2.
@@
    RE engineering suite of problems are set of tasks that are based on real-world applications in engineering, for instance ,rocket injector design and disc brake design. $d$ ranges from 3-7 variables and the number of objectives $m$ varies from 2 to 6. We use the same dataset as in prior work [37], which consists of 60,000 offline data points.
Baselines
We primarily compare our approach with two categories of baselines:
  1. 1.
@@
    We compare with ParetoFlow [40], a classifier guided generative model based on flow-matching [22]. Classifiers for guidance are trained surrogate models for each objective, followed by scalarization. It is important to note that ParetoFlow is our primary baseline since we are targeting inverse methods. In an inverse problem setting, the forward approaches introduced next are not applicable.
  2. 2.
@@
    Forward approaches using evolutionary algorithms: As suggested in prior work [37], a standard approach to offline MOO is to train a surrogate model for each objective and then use evolutionary algorithms such as NSGA-2 [10] to search over the design space. Although there are various ways to learn surrogate models, we compare with deep neural network (DNN)-based approaches, which are shown to perform best according to benchmarks [37]. The DNN approaches we compare with are: i) A Multi-Head Model: Uses multi-task learning [42] to train a joint surrogate for all objectives. Training techniques for this approach such as GradNorm [8] and PcGrad [39] are also compared. ii) Multiple Models: Maintain $m$ independent surrogate models, each making use of a single optimization technique, including COMs [33], ROMA [38], IOM [24], ICT [41], and Tri-mentoring [7].
Evaluation Metrics
@@
For each algorithm, we evaluate the convergence of solutions using the hypervolume metric [44], a standard metric in MOO for measuring the closeness of the proposed designs to the Pareto front. Hypervolume measures the volume of the objective space between a reference point and the objective vectors of the solution set, and does not require access to the true Pareto front.
@@
In addition to the hypervolume, we also measure the diversity of the obtained solutions using the $\Delta$-spread metric [10, 43]. The $\Delta$-spread measures the extent of the spread achieved in a computed Pareto front approximation [2]. It is important to consider the diversity of the obtained solutions, especially in the case of MOO wherein there is no single “best” design, but rather an entire set of solutions based on the Pareto front. In addition, in the case of offline optimization, the acquisition is single-shot. Therefore, solutions that are diverse and hence provide more coverage over the objective space are preferable. In this work, we provide the first effort to evaluate and benchmark the diversity of solutions obtained by different approaches in offline MOO. We evaluate all methods on five random seeds, and we compute the metrics using a budget of 256 designs.
@@
Figure 2: Plot of the samples from diffusion model (in green) on the ZDT2 task [45] at different timesteps of denoising. Top row shows results from our preference-guided diffusion, while the bottom row shows results from a purely unconditional model. Convergence of samples close to the Pareto front (in red) outside of the training data (blue) highlights the importance of preference guidance.
@@
Table 2: Hypervolume results of ZDT subtasks (part of the synthetic task) along with average rank of each method on the entire synthetic set of tasks. Each method is run for five random seeds and evaluated on 256 designs.
Method ZDT1 ZDT2 ZDT3 ZDT4 ZDT6 Avg. Rank
D (best) 4.17 4.67 5.15 5.45 4.61 8.43
MultiHead 4.8 $\pm$ 0.03 5.57 $\pm$ 0.07 5.58 $\pm$ 0.2 4.59 $\pm$ 0.26 4.78 $\pm$ 0.01 7.5
MultiHead - PcGrad 4.84 $\pm$ 0.01 5.55 $\pm$ 0.11 5.51 $\pm$ 0.03 3.68 $\pm$ 0.70 4.67 $\pm$ 0.1 6.08
MultiHead - GradNorm 4.63 $\pm$ 0.15 5.37 $\pm$ 0.17 5.54 $\pm$ 0.2 3.28 $\pm$ 0.9 3.81 $\pm$ 1.2 9.75
MultipleModels 4.81 $\pm$ 0.02 5.57 $\pm$ 0.07 5.48 $\pm$ 0.21 5.03 $\pm$ 0.19 4.78 $\pm$ 0.01 6.5
MultipleModels - COM 4.52 $\pm$ 0.02 4.99 $\pm$ 0.12 5.49 $\pm$ 0.07 5.10 $\pm$ 0.08 4.41 $\pm$ 0.21 7.83
MultipleModels - IOM 4.68 $\pm$ 0.12 5.45 $\pm$ 0.11 5.61 $\pm$ 0.06 4.99 $\pm$ 0.21 4.75 $\pm$ 0.01 5.58
MultipleModels - ICT 4.82 $\pm$ 0.01 5.58 $\pm$ 0.01 5.59 $\pm$ 0.06 4.63 $\pm$ 0.43 4.75 $\pm$ 0.01 6.67
MultipleModels - RoMA 4.84 $\pm$ 0.01 5.43 $\pm$ 0.35 5.89 $\pm$ 0.04 4.13 $\pm$ 0.11 1.71 $\pm$ 0.10 6.58
MultipleModels - TriMentoring 4.64 $\pm$ 0.10 5.22 $\pm$ 0.11 5.16 $\pm$ 0.04 5.12 $\pm$ 0.12 2.61 $\pm$ 0.01 8.33
ParetoFlow 4.23 $\pm$ 0.04 5.65 $\pm$ 0.11 5.29 $\pm$ 0.14 5.00 $\pm$ 0.22 4.48 $\pm$ 0.11 7.58
PGD-MOO + Data Pruning (Ours) 4.54 $\pm$ 0.08 5.21 $\pm$ 0.06 5.61 $\pm$ 0.06 5.06 $\pm$ 0.07 4.56 $\pm$ 0.14 5.08
PGD-MOO (Ours) 4.41 $\pm$ 0.08 5.33 $\pm$ 0.05 5.54 $\pm$ 0.10 5.02 $\pm$ 0.03 4.82 $\pm$ 0.01 4.42
@@ 5.1 Training Details
We parameterize the unconditional denoising model to be a multi-layer perceptron (MLP) with two 512-dimensional hidden layers, followed by a ReLU nonlinearity and layer normalization [21]. We also incorporate sinusoidal time embedding [35] for conditioning. We parameterize the preference model to be an MLP with three hidden layers, with first two hidden layers having the same number of units as the input, while the last hidden layer is having 512 units. Similar to denoising model, we also use ReLU nonlinearity followed by layer normalization and sinusoidal time embedding.
The denoising model is trained with AdamW optimizer [23] with learning rate of $45e-$ 4 for up to 200 epochs. The preference model is trained with Adam optimizer [15] with learning rate of $51e-$ 5 for up to 500 epochs. During sampling, we set the guidance weight $w$ to 10. For the preference model, we also experiment with pruning the training data to only contain the top $30\%$ of points, sorted according to their dominance. We refer to this method as PGD-MOO + Data Pruning in the results.
@@
Method RE21 RE22 RE25 RE32 RE35 RE37 RE41 RE61 Avg. Rank
D(best) 4.1 4.78 4.79 10.56 10.08 5.57 18.27 97.49 11.3
MultiHead-GradNorm 4.28 $\pm$ 0.39 4.7 $\pm$ 0.44 4.52 $\pm$ 0.5 10.54 $\pm$ 0.15 9.76 $\pm$ 1.3 5.67 $\pm$ 1.41 17.06 $\pm$ 3.82 108.01 $\pm$ 1.0 11.4
MultiHead-PcGrad 4.59 $\pm$ 0.01 4.73 $\pm$ 0.36 4.78 $\pm$ 0.14 10.63 $\pm$ 0.01 10.51 $\pm$ 0.05 6.68 $\pm$ 0.06 20.66 $\pm$ 0.1 108.54 $\pm$ 0.23 6.06
MultiHead 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.74 $\pm$ 0.2 10.6 $\pm$ 0.05 10.49 $\pm$ 0.07 6.67 $\pm$ 0.05 20.62 $\pm$ 0.11 108.92 $\pm$ 0.22 5.73
MultipleModels-COM 4.38 $\pm$ 0.09 4.84 $\pm$ 0.0 4.83 $\pm$ 0.01 10.64 $\pm$ 0.01 10.55 $\pm$ 0.02 6.35 $\pm$ 0.1 20.37 $\pm$ 0.06 107.99 $\pm$ 0.48 7.27
MultipleModels-ICT 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.64 $\pm$ 0.0 10.5 $\pm$ 0.01 6.73 $\pm$ 0.0 20.58 $\pm$ 0.04 108.68 $\pm$ 0.27 4.67
MultipleModels-IOM 4.58 $\pm$ 0.02 4.84 $\pm$ 0.0 4.83 $\pm$ 0.01 10.65 $\pm$ 0.0 10.57 $\pm$ 0.01 6.71 $\pm$ 0.02 20.66 $\pm$ 0.05 107.71 $\pm$ 0.5 4.6
MultipleModels-RoMA 4.57 $\pm$ 0.0 4.61 $\pm$ 0.51 4.83 $\pm$ 0.01 10.64 $\pm$ 0.0 10.53 $\pm$ 0.03 6.67 $\pm$ 0.02 20.39 $\pm$ 0.09 108.47 $\pm$ 0.28 7.67
MultipleModels-TriMentoring 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.62 $\pm$ 0.01 10.59 $\pm$ 0.0 6.73 $\pm$ 0.01 20.68 $\pm$ 0.04 108.61 $\pm$ 0.29 4.3
MultipleModels 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.63 $\pm$ 0.25 10.62 $\pm$ 0.02 10.55 $\pm$ 0.01 6.73 $\pm$ 0.03 20.77 $\pm$ 0.08 108.96 $\pm$ 0.06 3.67
ParetoFlow 4.2 $\pm$ 0.17 4.86 $\pm$ 0.01 - 10.61 $\pm$ 0.0 11.12 $\pm$ 0.02 6.55 $\pm$ 0.59 19.41 $\pm$ 0.92 107.1 $\pm$ 6.96 6.17
PGD-MOO + Data Pruning (Ours) 4.42 $\pm$ 0.04 4.83 $\pm$ 0.01 4.84 $\pm$ 0.0 10.64 $\pm$ 0.0 10.43 $\pm$ 0.04 5.99 $\pm$ 0.18 19.37 $\pm$ 0.15 103.04 $\pm$ 1.71 9.06
PGD-MOO (Ours) 4.46 $\pm$ 0.03 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.65 $\pm$ 0.0 10.32 $\pm$ 0.1 6.13 $\pm$ 0.12 19.31 $\pm$ 0.46 105.02 $\pm$ 1.14 7.67
Table 4: Average ranking of the $\Delta$-spread metric obtained by different algorithms on both synthetic and RE tasks. Detailed results are provided in Appendix A.

### v2

@@ 4.1 Preference Guided Diffusion
In this work, we explore an alternate guidance strategy that does not involve training surrogate models for every objective. Instead, we model the space of multiple-objectives through preference pairs, as defined by a Bradley-Terry model [bradley1952rank] of designs.
In this regard, we train a preference model that predicts the (log)- probability of whether a design Pareto dominates (Definition˜2.1) another design. During sampling of the designs, Given two designs $\bm{x}$} and $\hat{\bm{x}}$, we train a (time-conditioned) binary classifier that predicts $\log p_{\phi}({\bm{x}}\prec\hat{\bm{x}}\mid{\bm{x}},\hat{\bm{x}},t)$. We parameterize this distribution with a multi-layer perceptron (MLP) that takes in two inputs (designs) of size $2\times d$ and outputs the logit of the Bernoulli distribution predicting whether the first input Pareto dominates the second input.
@@
[…]
$d_{\text{CD}}({\bm{x}})=\sum_{i=1}^{m}\frac{y_{i}^{+}-y_{i}^{-}}{y_{i}^{\text{max}}-y_{i}^{\text{min}}},$ (7)
where $y_{i}^{+}$ and $y_{i}^{-}$ are the $i$th objective values of neighboring designs of $\bm{x}$} in the corresponding front sorted according to $i$th objective value. $y_{i}^{\text{max}}$ and $y_{i}^{\text{min}}$ are the maximum and minimum values of the objective $i$ in the current front. Crowding distance has been used as a secondary selection criterion in evolutionary algorithms such as NSGA-2 [deb2002fast] to maintain the diversity of solutions. Using crowding distance to create a binary label encourages the preference model to not only guide the diffusion model towards more Pareto-dominant regions, but also ensure that the designs that make up the Pareto front are diverse. For the denoising model, we train an unconditional diffusion model $\epsilon_{\theta}({\bm{x}}_{t},t)$ (§˜2.2) on the designs $\bm{x}$} in the dataset, similar to DDPM [ho2020denoising].
@@
[…]
3:for t = $T$ to 11 do
@@
With a trained denoising model $\epsilon_{\theta}({\bm{x}}_{t},t)$ and preference model $p_{\phi}({\bm{x}}\prec\hat{\bm{x}}\mid{\bm{x}},\hat{\bm{x}},t)$, we sample a new design by using classifier guidance (§˜2.3). We input both the denoised variable at the current timestep $\tilde{\bm{x}}_{t}$ as well as from the previous realization of denoising, i.e., $\tilde{\bm{x}}_{t+1}$ for the preference model to estimate $\nabla_{\tilde{\bm{x}}_{t}}p_{\phi}(\tilde{\bm{x}}_{t}\prec\tilde{\bm{x}}_{t+1}\mid\tilde{\bm{x}}_{t},\tilde{\bm{x}}_{t+1},t)$. At the beginning of the denoising procedure when ${\bm{x}}_{t+1}$ is not defined, we input the most dominant design in the dataset for preference comparison. The most dominant design in the dataset is chosen based on non-dominated sorting. The sampling procedure is summarized in Algorithm˜1. Intuitively, a preference model that generalizes well beyond its training data should guide the denoising process such that, at each step of denoising, the resulting sample $\tilde{\bm{x}}_{t}$ is in a more Pareto-dominant region.
With enough timesteps, the resulting denoised sample $\tilde{\bm{x}}_{0}$ will be close to the Pareto front. Moreover, the connection to Bradley-Terry model of preference is apparent: the gradient of the logit which indicates the probability of the first input dominating the second, carries information about the reward [rafailov2023direct]. Here, we use this to guide the diffusion model such that it can sample from regions closer to the Pareto front through the denoising process.
This approach does not require training surrogate models, thus providing a simple alternative approach to offline MOO. We find that the preference model generalizes well outside of the training data (see Fig.˜2), therefore providing guidance to the diffusion model to generate designs outside of the training data close to the Pareto front, while maintaining diversity in the samples.
@@ 5 Experiments
Our evaluation closely follows the benchmarking effort provided in prior work [xue2024offline]. We evaluate our approach on two sets of tasks: synthetic and real-world applications-based RE engineering suite [tanabe2020easy]. Each task consists of a dataset of 60k offline datapoints. As in [xue2024offline], we use 54k randomly chosen data points for training and the remaining for validation.
1. Synthetic set of tasks consist of 12 distinct tasks each with their own dataset. The objectives are analytic functions with tractable Pareto-fronts. This set of tasks has been widely used in MOO problems. Every task consists of 2-3 objectives with $d$ ranging from 10 to 30.
@@
2. RE engineering suite of problems are set of 12 distinct tasks, each with their own dataset. These tasks are based on real-world applications in engineering, for instance, rocket injector design and disc brake design. $d$ ranges from 3-7 variables and the number of objectives $m$ varies from 2 to 6.
3. MO-NAS are set of tasks are based on neural architecture search benchmarks from [dong2020bench, li2021hw, lu2023neural] which consists of multiple objectives like prediction error and hardware constraints like GPU latency. In particular, we use the datasets corresponding to C10/MOP(1-9) and IN1K/MOP(1-9) tasks provided by lu2023neural. As the design space is discrete, we operate on the continuous space of corresponding logits, as in prior work [xue2024offline, yuan2024paretoflow].
Details of individual tasks and their corresponding datasets are given in Appendix˜B.
Table 1: Average ranking of hypervolume obtained by different methods across synthetic, RE and MO-NAS tasks.
Method Synthetic RE MO-NAS
MultiHead 7.5 5.73 10.11
MultiHead - PcGrad 6.08 6.06 8.00
MultiHead - GradNorm 9.75 11.4 8.50
MultipleModels 6.5 3.67 9.39
MultipleModels - COM 7.83 7.27 2.72
MultipleModels - IOM 5.58 4.6 5.44
MultipleModels - ICT 6.67 4.67 6.61
MultipleModels - RoMA 6.58 7.67 7.44
MultipleModels - TriMentoring 8.33 4.3 8.11
ParetoFlow 7.58 6.17 4.67
PGD-MOO + Data Pruning (Ours) 5.08 9.06 2.14
PGD-MOO (Ours) 4.42 7.67 1.86
Baselines:
We compare our approach with two categories of baselines:
@@
1. We compare with ParetoFlow [yuan2024paretoflow], a classifier guided generative model based on flow-matching [lipman2022flow]. Classifiers for guidance are trained surrogate models for each objective, followed by scalarization. ParetoFlow is our primary baseline since our work is mainly concerned with studying inverse approaches for offline MOO.
@@
2. Forward approaches using evolutionary algorithms: As suggested in prior work [xue2024offline], a standard approach to offline MOO is to train a surrogate model for each objective and then use evolutionary algorithms such as NSGA-2 [deb2002fast] to search over the design space. Although there are various ways to learn surrogate models, we compare with deep neural network (DNN)-based approaches, which are shown to perform best according to benchmarks [xue2024offline]. The DNN approaches we compare with are: i) A Multi-Head Model: Uses multi-task learning [zhang2021survey] to train a joint surrogate for all objectives. Training techniques for this approach such as GradNorm [chen2018gradnorm] and PcGrad [yu2020gradient] are also compared. ii) Multiple Models: Maintain $m$ independent surrogate models, each making use of a single optimization technique, including COMs [trabucco2021conservative], ROMA [yu2021roma], IOM [qi2022data], ICT [yuan2024importance], and Tri-mentoring [chen2024parallel].
Other forward approaches like multi-objective Bayesian optimization algorithms are shown to perform worse than evolutionary algorithms on both synthetic and RE set of tasks [xue2024offline], hence we exclude them from comparisons in this work.
Evaluation Metrics:
@@
For each algorithm, we evaluate the convergence of solutions using the hypervolume metric [zitzler1998multiobjective], a standard metric in MOO for measuring the closeness of the proposed designs to the Pareto front. Hypervolume measures the volume of the objective space between a reference point and the objective vectors of the solution set, and does not require access to the true Pareto front. The reference point used for evaluation of hypervolume is taken from [xue2024offline](Appendix˜B).
@@
In addition to the hypervolume, we also measure the diversity of the obtained solutions using the $\Delta$-spread metric [deb2002fast, zhou2006combining]. The $\Delta$-spread measures the extent of the spread achieved in a computed Pareto front approximation [audet2021performance]. It is important to consider the diversity of the obtained solutions, especially in the case of MOO wherein there is no single “best” design, but rather an entire set of solutions based on the Pareto front. In addition, in the case of offline optimization, the acquisition is single-shot. Therefore, solutions that are diverse and hence provide more coverage over the objective space are preferable. In this work, we provide the first effort to evaluate and benchmark the diversity of solutions obtained by different approaches in offline MOO. Further details of the evaluation metrics and their computation is provided in §˜A.1.
We evaluate all methods on 5 random seeds, and compute the metrics using a budget of 256 designs.
@@
Figure 3: Plot of the samples from our preference-guided diffusion model (in green) on the ZDT2 task [zitzler2000comparison] at different timesteps of denoising. Convergence of samples close to the Pareto front (in red) outside of the training data (blue) highlights the importance of preference guidance.
@@
Table 3: Hypervolume results of ZDT subtasks (part of the synthetic task). Each method is run for five random seeds and evaluated on 256 designs.
Method ZDT1 ZDT2 ZDT3 ZDT4 ZDT6
D (best) 4.17 4.67 5.15 5.45 4.61
MultiHead 4.8 $\pm$ 0.03 5.57 $\pm$ 0.07 5.58 $\pm$ 0.2 4.59 $\pm$ 0.26 4.78 $\pm$ 0.01
MultiHead - PcGrad 4.84 $\pm$ 0.01 5.55 $\pm$ 0.11 5.51 $\pm$ 0.03 3.68 $\pm$ 0.70 4.67 $\pm$ 0.1
MultiHead - GradNorm 4.63 $\pm$ 0.15 5.37 $\pm$ 0.17 5.54 $\pm$ 0.2 3.28 $\pm$ 0.9 3.81 $\pm$ 1.2
MultipleModels 4.81 $\pm$ 0.02 5.57 $\pm$ 0.07 5.48 $\pm$ 0.21 5.03 $\pm$ 0.19 4.78 $\pm$ 0.01
MultipleModels - COM 4.52 $\pm$ 0.02 4.99 $\pm$ 0.12 5.49 $\pm$ 0.07 5.10 $\pm$ 0.08 4.41 $\pm$ 0.21
MultipleModels - IOM 4.68 $\pm$ 0.12 5.45 $\pm$ 0.11 5.61 $\pm$ 0.06 4.99 $\pm$ 0.21 4.75 $\pm$ 0.01
MultipleModels - ICT 4.82 $\pm$ 0.01 5.58 $\pm$ 0.01 5.59 $\pm$ 0.06 4.63 $\pm$ 0.43 4.75 $\pm$ 0.01
MultipleModels - RoMA 4.84 $\pm$ 0.01 5.43 $\pm$ 0.35 5.89 $\pm$ 0.04 4.13 $\pm$ 0.11 1.71 $\pm$ 0.10
MultipleModels - TriMentoring 4.64 $\pm$ 0.10 5.22 $\pm$ 0.11 5.16 $\pm$ 0.04 5.12 $\pm$ 0.12 2.61 $\pm$ 0.01
ParetoFlow 4.23 $\pm$ 0.04 5.65 $\pm$ 0.11 5.29 $\pm$ 0.14 5.00 $\pm$ 0.22 4.48 $\pm$ 0.11
PGD-MOO + Data Pruning (Ours) 4.54 $\pm$ 0.08 5.21 $\pm$ 0.06 5.61 $\pm$ 0.06 5.06 $\pm$ 0.07 4.56 $\pm$ 0.14
PGD-MOO (Ours) 4.41 $\pm$ 0.08 5.33 $\pm$ 0.05 5.54 $\pm$ 0.10 5.02 $\pm$ 0.03 4.82 $\pm$ 0.01
@@ 5.1 Training Details
We parameterize the unconditional denoising model to be a multi-layer perceptron (MLP) with two 512-dimensional hidden layers, followed by a ReLU nonlinearity and layer normalization [lei2016layer]. We also incorporate sinusoidal time embedding [vaswani2017attention] for conditioning. We parameterize the preference model to be an MLP with three hidden layers, with first two hidden layers having the same number of units as the input, while the last hidden layer is having 512 units. Similar to denoising model, we also use ReLU nonlinearity followed by layer normalization and sinusoidal time embedding.
The denoising model is trained with AdamW optimizer [loshchilov2017decoupled] with learning rate of 5e45e-4 for up to 200 epochs. Following ho2020denoising, we employ a linear noise schedule such that the noise $\beta_{t}$ grows linearly from 1e41e-4 to $0.02$. The preference model is trained with Adam optimizer [kingma2014adam] with learning rate of 1e51e-5 for up to 500 epochs. During sampling, we set the guidance weight $w$ to 10. For the preference model, we also experiment with pruning the training data to only contain the top $30\%$ of points, sorted according to their dominance. We refer to this method as PGD-MOO + Data Pruning in the results.
@@
Method RE21 RE22 RE25 RE32 RE35 RE37 RE41 RE61
D(best) 4.1 4.78 4.79 10.56 10.08 5.57 18.27 97.49
MultiHead-GradNorm 4.28 $\pm$ 0.39 4.7 $\pm$ 0.44 4.52 $\pm$ 0.5 10.54 $\pm$ 0.15 9.76 $\pm$ 1.3 5.67 $\pm$ 1.41 17.06 $\pm$ 3.82 108.01 $\pm$ 1.0
MultiHead-PcGrad 4.59 $\pm$ 0.01 4.73 $\pm$ 0.36 4.78 $\pm$ 0.14 10.63 $\pm$ 0.01 10.51 $\pm$ 0.05 6.68 $\pm$ 0.06 20.66 $\pm$ 0.1 108.54 $\pm$ 0.23
MultiHead 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.74 $\pm$ 0.2 10.6 $\pm$ 0.05 10.49 $\pm$ 0.07 6.67 $\pm$ 0.05 20.62 $\pm$ 0.11 108.92 $\pm$ 0.22
MultipleModels-COM 4.38 $\pm$ 0.09 4.84 $\pm$ 0.0 4.83 $\pm$ 0.01 10.64 $\pm$ 0.01 10.55 $\pm$ 0.02 6.35 $\pm$ 0.1 20.37 $\pm$ 0.06 107.99 $\pm$ 0.48
MultipleModels-ICT 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.64 $\pm$ 0.0 10.5 $\pm$ 0.01 6.73 $\pm$ 0.0 20.58 $\pm$ 0.04 108.68 $\pm$ 0.27
MultipleModels-IOM 4.58 $\pm$ 0.02 4.84 $\pm$ 0.0 4.83 $\pm$ 0.01 10.65 $\pm$ 0.0 10.57 $\pm$ 0.01 6.71 $\pm$ 0.02 20.66 $\pm$ 0.05 107.71 $\pm$ 0.5
MultipleModels-RoMA 4.57 $\pm$ 0.0 4.61 $\pm$ 0.51 4.83 $\pm$ 0.01 10.64 $\pm$ 0.0 10.53 $\pm$ 0.03 6.67 $\pm$ 0.02 20.39 $\pm$ 0.09 108.47 $\pm$ 0.28
MultipleModels-TriMentoring 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.62 $\pm$ 0.01 10.59 $\pm$ 0.0 6.73 $\pm$ 0.01 20.68 $\pm$ 0.04 108.61 $\pm$ 0.29
MultipleModels 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.63 $\pm$ 0.25 10.62 $\pm$ 0.02 10.55 $\pm$ 0.01 6.73 $\pm$ 0.03 20.77 $\pm$ 0.08 108.96 $\pm$ 0.06
ParetoFlow 4.2 $\pm$ 0.17 4.86 $\pm$ 0.01 - 10.61 $\pm$ 0.0 11.12 $\pm$ 0.02 6.55 $\pm$ 0.59 19.41 $\pm$ 0.92 107.1 $\pm$ 6.96
PGD-MOO + Data Pruning (Ours) 4.42 $\pm$ 0.04 4.83 $\pm$ 0.01 4.84 $\pm$ 0.0 10.64 $\pm$ 0.0 10.43 $\pm$ 0.04 5.99 $\pm$ 0.18 19.37 $\pm$ 0.15 103.04 $\pm$ 1.71
PGD-MOO (Ours) 4.46 $\pm$ 0.03 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.65 $\pm$ 0.0 10.32 $\pm$ 0.1 6.13 $\pm$ 0.12 19.31 $\pm$ 0.46 105.02 $\pm$ 1.14
5.2 Results
Evaluation of Convergence. We provide detailed results of hypervolume for various baselines and our approach on the synthetic task (Tables˜2 and 3). We find that our approach performs competitively with respect to baselines. Preference-guided diffusion performs on average better than ParetoFlow, another generative model-based approach using guidance. This shows the benefits of having a preference model as a classifier for guidance. Vizualization of the final sampled designs from our approach is provided in §˜D.2. Overall, our method performs better than other baselines, which learn surrogate models and use evolutionary algorithms in the synthetic task setting (Fig.˜3). In addition, we also find that our method performs competitively in the RE engineering suite (Tables˜4 and 22) and the MO-NAS tasks (Table˜1). In problems with higher number of objectives, we find that our approach is slightly worse compared to the baselines in terms of hypervolume. However, we note that our approach is much simpler to train in these settings, while still achieving diverse solutions (discussed further below).
Evaluation of Diversity. Average ranking in terms of performance of the $\Delta$-spread metric for all algorithms (Table˜5) shows that our approach gives more diverse solutions than all the other baselines including in the RE setting. These results highlight the importance of having a diversity constraint in the training procedure for the classifier through the data selection procedure (§˜D.3).
Across experiments, we find that our approach has competitive performance in terms of hypervolume (convergence) while being better in terms of the $\Delta$-spread metric (diversity) than the baselines.
Table 5: Average ranking of the $\Delta$-spread metric obtained by different algorithms on both synthetic and RE tasks. Detailed results are provided in Appendix˜D.